If the perimeter of a rectangle is 60 inches. If the width of the rectangle is 12 inches, then what is the length of the rectangle?
Answer by checkley77(12844)

P=2L+2W
60=2L+2*12
60=2L+24
2L=60-24
2L=36
L=36/2
L=18 ANS. FOR THE LENGTH.
PROOF:
60=2*16+2*12
60=36+24
60=60